OpenStudy (anonymous):

Okay I need to solve this equation and be able to come up with a linear equation and if someone could do the steps awesome y=9x-4 I have gotten responses but when I use -1,0,1 it does not come out right for my graph

OpenStudy (gorv):

this is linear equation

OpenStudy (anonymous):

Okay what do I do to solve it so I get three points to graph

OpenStudy (anonymous):

it's already solved, it's a linear equation

OpenStudy (gorv):

take any point ..hmmmm take x=0

OpenStudy (anonymous):

But my video shows you need to break it down

OpenStudy (anonymous):

first point: set x = 0, you get y = -4 -> (0, -4) Second point: set y=0, you get x = 4/9 -> (4/9, 0) Third point: set x = 1, you get y = 9 - 4 = 5 -> (1, 5)

OpenStudy (gorv):

at x=0 find y y=9*0-4 y=-4 (0,-4)

OpenStudy (campbell_st):

use a simple table of values pick the x values, such as x: 0 : 1 : 2 ----------------------------------- y: 9(0) - 4 = -4 : 9(1) - 4 = : 9(2) - 4 = substitute the x values to get the corresponding y values then plot the ordered pairs... (0, -4) etc.. hope it makes sense

OpenStudy (anonymous):

It does make sense but I am being told to use a negative number like -1,0,1 and when I do that I come up with 0,-4, 9,5, -9,-13 and it doesn't work on my graph I feel really stupid right now and am beyond frustrated

OpenStudy (campbell_st):

well it's the same thing x: -1 : 0 : 1 --------------------------------------- y: 9(-1) - 4 = -13 :9(0) - 4 = : 9(1) - 4 = you can choose any 3 points you like. The reasoning for 3 points is that is you make a mistake on 1 point then the 3 points won't form a straight line. I always choose x = 0 as its the y- intercept.

OpenStudy (anonymous):

When I graph it on my online homework it doesn't graph correctly so I don't know what I am doing wrong I feel dumb